OpenStudy (anonymous):

Can Someone Help Me? Please? Metals, Fans , Anything...! Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years

OpenStudy (anonymous):

Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

Well, an exponential growth would be of the form: \[A = A_oe^{Bt}\] Where A is what you have at time t, Ao is what you started with, B is a growth parameter that controls how fast the population would grow, and t is time.

OpenStudy (anonymous):

it doesnt tell me anything

OpenStudy (anonymous):

Everything you need is there. Ao is what you start with. What is the population that you start with?

OpenStudy (anonymous):

i doesnt tell me

OpenStudy (anonymous):

Right. But we can say the "initial population," which I believe is what you're calling the "principle population."

OpenStudy (anonymous):

meaning?

OpenStudy (anonymous):

the exponential term (e^Bt) is what controls the rate of growth. B in particular. The larger the value of B, the faster the population will grow.

OpenStudy (anonymous):

ugh im confused /:

OpenStudy (anonymous):

Have a calculator?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Let's play around with this equation a bit. First, let's look at the exponential term. First, let's choose B=1. We'll look at times of t =1, t=5, and t=10. So, t = 1, 5, 10 For t=1, we get e^1 For t = 5, we get e^5 For t = 10, we get e^10 Next, let's choose B=10. We'll still look at the same values for t. Bt = 10, 50, 100 For t = 1, we get Bt = 10. So: e^10 For t=5, we get Bt = 50. So: e^50 For t=10, we get Bt = 100. So: e^100 If you evaluate each of these, you'll see that the larger B is, the faster the exponential grows (keep in mind that we used exactly the same times for each of these)

OpenStudy (anonymous):

@sleepyjess could you help me? im not understanding him... please

OpenStudy (anonymous):

I think I did this, but hold on, I have to finish this test first. :)

OpenStudy (anonymous):

thank you !! your my life saver <3

OpenStudy (anonymous):

Ur welcome! Wait for me, I'll message it to you. :)

OpenStudy (anonymous):

thank you x a million

OpenStudy (anonymous):

Ur welcome! Haha. :)